# H-FA: A Hybrid Floating-Point and Logarithmic Approach to Hardware Accelerated FlashAttention

### v1

@@ VI-D Comparison with other state-of-the-art designs
For completeness, we compare the proposed design with other attention accelerators that do not necessarily adopt the fused computation paradigm introduced by the FlashAttention algorithm. For example, state-of-the-art designs used in our comparison rely on generic systolic architectures or multiple vector units for matrix multiplication, while employing custom-designed units to compute non-linear activation functions such as softmax [16, 9, 12, 15, 11, 19, 14, 18]. Among them, only COSA plus [14] and TSAcc [18] target specifically attention acceleration.
@@
The proposed designs, which apply logarithmic computation to carefully selected parts of the FlashAttention algorithm, deliver comparable energy and area efficiencies with only a small overhead in area and power. Compared to attention-specific accelerators such as COSA plus [14] and TSAcc [18], which also uses similar floating-point datatypes, H-FA-1-4 and H-FA-4-4 achieve significantly higher area and energy efficiencies. COSA plus [14] employs large systolic arrays and can offer significant throughput. However, its energy efficiency is reduced due to its FPGA implementation.

### v2

@@ VI-D Comparison with other state-of-the-art designs
For completeness, we compare the proposed design with other attention accelerators that do not necessarily adopt the fused computation paradigm introduced by the FlashAttention algorithm. For example, state-of-the-art designs used in our comparison rely on generic systolic architectures or multiple vector units for matrix multiplication, while employing custom-designed units to compute non-linear activation functions such as softmax [chen, keller, kim, moon, mecla, fact, cosa, tsacc]. Among them, only COSA plus [cosa] and TSAcc [tsacc] target specifically attention acceleration.
@@
The proposed designs, which apply logarithmic computation to carefully selected parts of the FlashAttention algorithm, deliver comparable energy and area efficiencies with only a small overhead in area and power. Compared to attention-specific accelerators such as COSA plus [cosa] and TSAcc [tsacc], which also uses similar floating-point datatypes, H-FA-1-4 and H-FA-4-4 achieve significantly higher area and energy efficiencies. COSA plus [cosa] employs large systolic arrays and can offer significant throughput. However, its energy efficiency is reduced due to its FPGA implementation.